OpenStudy (anonymous):

Factor completely 12x^2 - 11x + 2 and x^2 - 49

hero (hero):

To factor a trinomial of the form \(ax^2 + bx + c\), you perform the following 1. Find two numbers \(m\) and \(n\) that add to get \(a \times c\) yet also add to get \(b\). 2. Replace \(b\) with \(m + n\) 3. Factor by grouping To factor a binomial square of the form a^2 - b^2, you simply use the formula \(a^2 - b^2 = (a + b)(a - b)\)

OpenStudy (anonymous):

thanks hero.

hero (hero):

So to factor \(12x^2 - 11x + 2\), 1. \(m \times n = 24\) \(m + n = -11\) @melf1717, can you think of two numbers that multiply to get 24, yet add to get -11?

OpenStudy (anonymous):

for ax^2+bx+c=0 --> (x+x1)(x+x2)=0 when x1 and x2 = \[ = \frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

\[=\frac{ 11\pm \sqrt{121-4(12)(2)} }{ 2(12) }=\frac{ 11\pm5 }{ 24 }\]

OpenStudy (anonymous):

x1=16/24 and x2=6/24 then (x1+2/3)(x+1/4)

hero (hero):

I get -8 and -3 for the two numbers and then from that we know that -11 = -(3 + 8) So the factorization becomes 12x^2 - (3 + 8)x + 2 = 12x^2 - 3x - 8x + 2 = 3x( 4x - 1) - 2(4x - 1) = (4x - 1)(3x - 2)